Solve sin(20) NIE for 0 <o<2m. Give your answer in radians.
Answer #1

= WIS sol Sin (201 sin (20) = sin(7) sim (90+2) = sin ( 201 + $7%) = 201 - 50 - 407 = { 1 2 So 20 2. ㅈ so 1301 I 170 12 12